# 3DGAA: Realistic and Robust 3D Gaussian-based Adversarial Attack for Autonomous Driving

### v2

@@ C.3 Optimization Efficiency
The efficiency of 3DGAA is a result of its computational design and strategic optimization approach. Unlike conventional adversarial attack frameworks that rely on texture-based deformations, 3DGAA operates on 3DGS representations, significantly reducing computational overhead. By leveraging the Gaussian-based parameterization, the attack directly optimizes object attributes without needing explicit surface reconstruction, leading to faster convergence and lower memory consumption. Usually, we complete an entire generative-adversarial process in less than one minute.
Additionally, the optimization process in 3DGAA follows a structured loss-balancing strategy. The adversarial loss and shape preservation loss are dynamically weighted, ensuring that the optimization process remains stable while effectively guiding the object toward an adversarial state. This selective gradient update mechanism minimizes unnecessary modifications, accelerating convergence without sacrificing attack strength.
Furthermore, the physical augmentation module is designed to enhance optimization stability by progressively introducing perturbations. This staged approach prevents destabilization during training, contributing to the overall computational efficiency of 3DGAA. Combining these design choices ensures that 3DGAA achieves high attack performance while maintaining physical reality in real-world adversarial scenarios.
C.4 Ethical Considerations

### v3

@@ C.2 Generalization to Non-Vehicle Objects
Table C.2: Efficiency profiling under a shared protocol (same data, views, and early-stopping rule). Times are wall-clock from process start to end; Total includes 3DGS generation, adversarial training, and data I/O/misc.
[…]
(Model) (MB) (s) (epochs/iters) (s) (s) (s) (s) (%)
3DGAA (Full) RTX 4090 23256 $9.24\pm 0.26$ $27$ $0.37$ $10.02\pm 0.22$ $32.26\pm 0.01$ $51.52$ $24378$ $98.22$
$\textbf{3DGAA}^{a}$ RTX 4090 19862 $9.24\pm 0.26$ $283$ $0.26$ $74.82\pm 3.20$ $32.26\pm 0.01$ $116.32$ $24378$ $80.25$
$\textbf{3DGAA}^{g}$ RTX 4090 21196 $9.24\pm 0.26$ $24$ $0.31$ $7.42\pm 0.15$ $32.26\pm 0.01$ $48.92$ $24378$ $75.05$
FCA [40] RTX 4090 23256 200 12.68 $2536\pm 96$ $92.52\pm 12.18$ 2629
TT3D [16] RTX 4090 23256 200 21.80 $4360\pm 74$ $166.29\pm 48.40$ 4529
@@ C.3 Optimization Efficiency
Protocol and measurements. We profile three modes (Full, $\text{3DGAA}^{a}$, $\text{3DGAA}^{g}$) under a shared setup (same data, views, early-stopping rule) on a single RTX 4090 with fixed driver/CUDA/PyTorch. Wall-clock time is reported from process start to end and decomposed into 3DGS Gen (initial Gaussian generation), Adv. Time (adversarial optimization), and Data/Misc (I/O, logging, warm-up). Adv. Iters denotes the final iteration count under the identical stopping criterion; Time/Iter is the mean per-iteration duration measured after device synchronization. Peak Mem is the maximum GPU memory in MB. We further report $G$ (the number of Gaussians after TP+SD; determined by the backbone and inputs) and $\Delta G$ (the percentage of Gaussians updated during adversarial optimization). For context, we also include two non-3DGS baselines: FCA [40] (fixed-mesh/texture pipeline) and TT3D [16] (NeRF-based). For these methods, 3DGS Gen is not applicable (fixed mesh or radiance-field training), so we report their adversarial optimization and I/O times under the same capture/view protocol. See Tab. C.2 for consolidated numbers.
$\text{3DGAA}^{a}$ attains the lowest Peak Mem (19,862 MB) and the fastest Time/Iter (0.26 s), but requires more iterations and thus a longer Adv. Time (74.82 s) and Total (116.32 s), which is verified in Fig. 5(a). In contrast, jointly optimizing appearance and geometry (Full) reaches the target earlier (27 iters; 10.02 s adversarial time) despite a higher per-iteration cost (0.37 s) and memory (23,256 MB), yielding a shorter end-to-end time (51.52 s). $\text{3DGAA}^{g}$ sits in between on memory (21,196 MB) and per-iteration time (0.31 s), converging in 24 iterations for a 48.92 s total. When contrasted with FCA and TT3D, the 3DGAA family operates at a markedly different efficiency scale: FCA and TT3D exhibit per-iteration costs in the order of tens of seconds (12.68 s / 21.80 s) and hours-level totals (2,629 s / 4,529 s), even without a 3DGS generation stage, whereas all 3DGAA modes finish within seconds-to-minutes under the same protocol. Overall, these findings substantiate the practicality of 3DGAA; detailed runtime numbers are provided here to illustrate the efficiency of 3DGAA, while the main paper focuses on realism and attack efficacy.
C.4 Direct Optimization on Segmentation
[…]
Table C.3: Segmentation model (DeepLabv3-R50) optimizing.
Method Vanilla Det $\rightarrow$ Seg Seg
LCR $\uparrow$ 0.000 0.628 1.144
[…]
C.5 Ethical Considerations